find Sn fort he given arithmetic series.. a1=22 d=-8 n=21

Answer by jhadipak(3)

Solution: we know that Sn = (n/2)(2a + (n - 1)d)
where n = last term
a = first term and d = difference between two terms.
Datas We have given in the problem as a1=22 d=-8 n=21
where a1 = 1st term
plugging all the given values will result
Sn = (n/2)(2a + (n - 1)d)
= (21/2)(2(22) + (21 - 1)(-8))
= (21/2)(44 + (20)(-8))
= (21/2)(44 - 160)
= (21/2)(-116)
= 21(-58)
= -1218
As we can see that the difference of the number given is negative so the last term must be negative.
Hope this will help you!!